OpenStudy (anonymous):

(ALGEBRA II) PLEASE HELP! 12a² - 3ab + 2b ; a= -5, b= 4

OpenStudy (lgbasallote):

sounds like substitution 12(-5)^2 - 3(-5)(4) + 2(4) what do you get as answer?

OpenStudy (anonymous):

I got 3,668 but my dad and I are debating over the answer, but he said that when he did 12a² he got 300 and that's the part we're really confused about

OpenStudy (anonymous):

But 3,668 was my overall answer for the whole problem. The 300 answer from him was only for the 12a² part

OpenStudy (anonymous):

i got 368...

OpenStudy (lgbasallote):

maybe this can help http://www.wolframalpha.com/input/?i=12%28-5%29%5E2+-+3%28-5%29%284%29+%2B+2%284%29

OpenStudy (anonymous):

@montanagolfer When you did it, did you do the (-5)² first?? and then multiply it by 12??

OpenStudy (anonymous):

use the wolframalpha if your never sure...its magic for helping on math problems

OpenStudy (anonymous):

@lgbasallote Thank you very much :)

OpenStudy (lgbasallote):

12(-5)^2 = 12 x 25 = 300 3(-5)(4) = -15 x 4 = -60 2(4) = 8 300 -(-60) + 8 = 300 + 60 + 8 = 368

OpenStudy (anonymous):

@lgbasallote @montanagolfer Ohhh I see it now, thanks :)

OpenStudy (lgbasallote):

wonderful